OpenStudy (anonymous):

By changing to polar coordinates, evaluate the double integral∬(x^(2)+y(2))^(5/2)dxdy where D is the disk x^(2)+y^(2)≤16.

zepdrix (zepdrix):

Mmmm ok so this is the region we're integrating over, yes?

zepdrix (zepdrix):

So in polar our function will be \(\Large f(r,\;\theta)\) Let's use the picture to try and change the limits of integration. So what will our boundaries be for r and theta? Lemme know if that's too confusing.

OpenStudy (anonymous):

theta is 0 to 2pi, r is 0 to 4?

zepdrix (zepdrix):

Good good good. Then we'll use these to change from cartesian to polar:\[\Large x=r \cos \theta, \qquad\qquad\qquad y=r \sin \theta\]

zepdrix (zepdrix):

When we change our differentials, we need to remember that an extra factor of r shows up.\[\Large dy\;dx \qquad\to\qquad r\;dr\;d\theta\]

OpenStudy (anonymous):

so the new integral is \[\int\limits_{0}^{2\pi}\int\limits_{0}^{4}r^5rdrd \theta\]

zepdrix (zepdrix):

ah yes! very good :)

OpenStudy (anonymous):

Thank you very much! You have been very helpful. Could you help me with another one?

zepdrix (zepdrix):

sure

OpenStudy (anonymous):

Calculate the integral over the given region by changing to polar coordinates: \[f(x,y)=\left| 16xy \right|, x^2+y^2\le 81\]

zepdrix (zepdrix):

Hmm ok lemme think a sec

zepdrix (zepdrix):

Hmm so we have nice easy conversion of the boundaries again, right?

OpenStudy (anonymous):

Yes, r spans from 0 to 9. I am unsure of the boundaries for theta though. Does the absolute value affect them? Or are they still 0 to 2pi?

zepdrix (zepdrix):

We still want to cover this entire region. So theta will still be 0 to 2pi.